Monthly Expenses (S$) Frequency
0 and less than 150 5
150 and less than 300 8
300 and less than 450 15
450 and less than 600 9
600 and less than 750 17
750 and less than 900 6
Total 60

If 80% of the monthly expenses is below $600, then it indicates that the XYZ has not been overspending. From the above grouped data, would you say that XYZ has been overspending? show your calculation.
1
Expert's answer
2019-05-27T14:30:19-0400
"Monthly \ Expenses \ (S \$) \ \ \ \ Frequency""\ \ 0\leq S<150 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 5""150\leq S<300 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 8""300\leq S<450 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 15""450\leq S<600 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 9""600\leq S<750 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 17""750\leq S<900 \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ 6"

Total 60 


"S<600: \ 5+8+15+9=37"

"80 \% \ of \ 60=48>37"

We would say that XYZ has been overspending.